What is x in this equation: 9(x+1)=2(3x+8)?
Answer by TimothyLamb(4379)

Use algebraic simplification, like so:
---
9(x+1) = 2(3x+8)
9x + 9 = 6x + 16
9x - 6x = 16 - 9
3x = 7
x = 7/3
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php